Shell is experiencing rapid growth. Earnings and dividends are expected to grow at a rate of 15% during the next 2 years, at 13% the following year, and at a constant rate of 6% during Year 4 and thereafter. Its last dividend was $1.15, and its required rate of return is 12%.

f) Calculate the dividend and capital gains yields for Years 1, 2, and 3.

Dividend Yield Year 1 =  %

Capital Gains Yield Year 1 =  %

Dividend Yield Year 2 =  %

Capital Gains Yield Year 2 =  %

Dividend Yield Year 3 =  %

Capital Gains Yield Year 3 =  %
Answer #1

Dividend Year 1=1.15*(1+15%)
Dividend Year 2=1.15*(1+15%)^2
Dividend year 3=1.15*(1+15%)^2*(1+13%)
Dividend year 4 =1.15*(1+15%)^2*(1+13%)*(1+6%)

Price in Year 3=Dividend year 4/(Rate-Growth) =1.15*(1+15%)^2*(1+13%)*(1+6%)/(12%-6%) =30.3617
Price In year 2= (Dividend Year 3 +Price in year 3)/(1+r) =(1.15*(1+15%)^2*(1+13%)+30.3617)/(1+12%) =28.6431
Price in year 1 =(Dividend Year 2 +Price in year 2)/(1+r) =(1.15*(1+15%)^2+28.6431)/(1+12%) =26.9321
Price in year 0 =(Dividend Year 1 +Price in year 1)/(1+r) =(1.15*(1+15%)^1+26.9321)/(1+12%) =25.2273

Dividend Yield Year 1=Dividend 1/Price 0 =1.15*(1+15%)/25.2273=5.24%
Capital Gain Year 1 =Required Rate-Dividend Yield =12%-5.24% =6.76%
Dividend Yield Year 2=Dividend 2/Price 1 =1.15*(1+15%)^2/26.9321=5.65%
Capital Gain Year 2 =Required Rate-Dividend Yield =12%-5.65% =6.35%
Dividend Yield Year 3=Dividend 2/Price 2 =1.15*(1+15%)^2*(1+6%)/28.6431=5.63%
Capital Gain Year 3 =Required Rate-Dividend Yield =12%-5.63% =6.37%

​​​​​​​

> helpful

paul chairuka Wed, Nov 17, 2021 12:28 PM